OpenStudy (anonymous):

Integral of x*sin(x^2) , using u substitution if necessary?

OpenStudy (luigi0210):

Yes, you could use a U sub

OpenStudy (luigi0210):

is there anything there that is a derivative of another?

OpenStudy (anonymous):

i kind of cheated and googled it, so, looks like the best bet is to set u = x^2, thereby, we can set x = 1/2 * du

OpenStudy (luigi0210):

\(\large u=x^2 \) and \(\large du=\frac{1}{2}x du \) Then yes

OpenStudy (luigi0210):

That should be dx at the end not du, sorry

OpenStudy (amistre64):

and the derivative of x^2 is not 1/2 x

OpenStudy (luigi0210):

Yea, meant to put that outside near du, this Is embarrassing \(\Large du=2x dx\ )

OpenStudy (anonymous):

think I got it from there, will post the larger problem for completeness, thanks for the help